OpenStudy (anonymous):

the table shows values for a function x|1|2|3 | 4| y|4|9|16|25| which equation represents this function A. y=x^2 B. y=x^2+1 C. y= (x - 1)^2 D. y= (x + 1)^2

OpenStudy (welshfella):

plug in values of x and y and see if they fit

OpenStudy (anonymous):

Just plug it in, Using rule D (1+1)^2 = 4 (2+1)^2= 9 D is ur answer

OpenStudy (welshfella):

for example y = x^2 y = 4 and x = 1 y = 4 x^2 = 1^2 = 1 - not equal so ist not A